A man has P1000 invested part of 5% and his remainder at 3% simple interest. The total income per year from these investment is P44. How much does he invested at the higher rate? P= Peso
Answer by rushdi1987(3)

I = PNR where I = interest , P = principal, N = time, R = rate of interest
Let he invested peso x at 5% interest and the rest (1000-x) @ 3% interest.
For 5% rate interest will be x*5/100*1 or x/20
For 3% rate interest will be 1000-x*3/100*1 or 3(1000-x)/100
According to the qus
x/20 + 3(1000-x)/100 = 44
or, 5x + 3(1000-x) = 44 * 100
or, 5x + 3000 - 3x = 4400
or, 2x = 4400-3000
or, 2x = 1400
or, x = 700
he invested p 700 at higher rate